How many ounces of pure zinc must be added to 20 ounces of brass which is 30% zinc and 70% copper to produce brass that is 40% zinc?
Answer by mananth(16949)

percent ---------------- quantity
30 ---------------- 20 brass
100 ---------------- x Zinc
40 ---------------- 20+x Mixture

30*20+100*x=40(20+x)

600+100 x=800+40x

100x-40x=800-600

60x=200

/60
x=3.33 ounces of pure Zinc